# A2Seek: Towards Reasoning-Centric Benchmark for Aerial Anomaly Understanding

### v1

@@ Abstract
While unmanned aerial vehicles (UAVs) offer wide-area, high-altitude coverage for anomaly detection, they face challenges such as dynamic viewpoints, scale variations, and complex scenes. Existing datasets and methods, mainly designed for fixed ground-level views, struggle to adapt to these conditions, leading to significant performance drops in drone-view scenarios. To bridge this gap, we introduce A2Seek (Aerial Anomaly Seek), a large-scale, reasoning-centric benchmark dataset for aerial anomaly understanding. This dataset covers various scenarios and environmental conditions, providing high-resolution real-world aerial videos with detailed annotations, including anomaly categories, frame-level timestamps, region-level bounding boxes, and natural language explanations for causal reasoning. Building on this dataset, we propose A2Seek-R1, a novel reasoning framework that generalizes R1-style strategies to aerial anomaly understanding, enabling a deeper understanding of “Where” anomalies occur and “Why” they happen in aerial frames. To this end, A2Seek-R1 first employs a graph-of-thought (GoT)-guided supervised fine-tuning approach to activate the model’s latent reasoning capabilities on A2Seek. Then, we introduce Aerial Group Relative Policy Optimization (A-GRPO) to design rule-based reward functions tailored to aerial scenarios. Furthermore, we propose a novel “seeking” mechanism that simulates UAV flight behavior by directing the model’s attention to informative regions. Extensive experiments demonstrate that A2Seek-R1 achieves up to a 22.04% improvement in AP for prediction accuracy and a 13.9% gain in mIoU for anomaly localization, exhibiting strong generalization across complex environments and out-of-distribution scenarios. Our dataset and code will be released at https://hayneyday.github.io/A2Seek/.
@@ 1 Introduction
Traditional anomaly detection lu2013abnormal ; liu2018ano_pred ; cao2023new relies on fixed-view cameras and primarily focuses on anomaly classification, offering limited semantic interpretation. Their static perspectives and narrow fields of view significantly limit their effectiveness in monitoring large and dynamic environments zhu2024advancing . With the rapid advancement of unmanned aerial vehicle (UAV) technology, aerial surveillance has emerged as a powerful paradigm for wide-area anomaly detection. Drone-view footage introduces frequent viewpoint shifts, scale changes, complex backgrounds and occlusions, as well as environmental disturbances (lighting, weather, moving shadows) cao2021visdrone ; du2018unmanned . Crucially, anomalous regions in aerial scenes are often subtle, spatially sparse, and occupy only a small portion of the field of view, making them difficult to perceive. Even when alarms are triggered, human observers often struggle to localize these subtle, spatially sparse events. Moreover, traditional models rely on stable spatial priors invalidated when a drone’s pose constantly changes. Consequently, ground-view approaches fail to generalize to aerial data, exhibiting poor robustness to dynamic scenes and variable spatial distributions. As Figure 1 illustrates, anomaly detection in drone-view footage requires precise spatial localization (“Where is the anomaly?”) and semantic interpretation (“Why is it an anomaly?”). Drone-captured videos often span large, dynamic environments, where subtle anomalies can be easily overlooked or misinterpreted without robust contextual reasoning. These factors demand models capable of adaptively focusing on critical regions and abstracting high-level reasoning to explain anomalies. In the following, we distill these into two core problems critical for robust aerial anomaly understanding.
@@
CHALLENGE 1: Spatial Localization in Wide-Area, Dynamic Scenes. Traditional ground-based anomaly detection typically relies on fixed-view cameras, where anomalies are often prominent and relatively easy to identify. However, aerial anomaly detection is fundamentally different. Drone-view videos cover expansive and dynamic scenes with complex backgrounds and frequent motion. Anomalies are often small, sparse, and easily obscured by cluttered surroundings, making them difficult to detect even when anomalies are flagged (e.g., a person falling in Figure 1a). Therefore, aerial anomaly detection requires not only determining whether an event is anomalous but also providing accurate localization of the anomalous objects for supporting anomaly judgment.
@@
Building on this benchmark, we propose A2Seek-R1, a novel reinforcement fine-tuning framework designed to enhance the reasoning capabilities of models for aerial anomaly understanding. A2Seek-R1 first employs a graph-of-thought (GoT)-guided supervised fine-tuning (SFT) approach, which activates the model’s latent reasoning capabilities by leveraging structured reasoning annotations in the A2Seek dataset. These annotations consist of optional stages, including trigger, diagnosis, reasoning, reflection, and seeking, effectively guiding the model to handle anomalies of varying complexity in a progressive manner. Among them, seeking is set as a potential region of interest for the model in video frames with insufficient information such as blurry or occluded images, thus achieving a new type of seeking mechanism that simulates the flight behavior of unmanned aerial vehicles, enabling the model to dynamically focus on specific regions of interest. Second, Second, it introduces a tailored extension of Group Relative Policy Optimization (GRPO), termed A-GRPO, specifically designed for aerial anomaly understanding. A-GRPO extends the original accuracy and format function rewards by incorporating localization and seeking rewards. Localization rewards enhance the model’s spatial understanding of anomaly regions, while seeking rewards focus on aligning the model’s predictions with human annotations of anomaly candidate areas, ensuring the extraction of valuable spatial information for better understanding. Additionally, to address the diverse perspectives of drones, a length reward function is introduced to encourage concise responses in simple scenarios and allocate more computational effort to complex situations. By combining these components, A2Seek-R1 achieves precise spatial localization and robust reasoning for aerial anomalies, setting a new benchmark for anomaly undersatnding in complex, real-world environments.
@@ 2 Related Work
Video Anomaly Detection. Early efforts focused on single-scene datasets li2013anomaly ; lu2013abnormal using fixed-view RGB cameras for pedestrian anomaly detection. Later datasets liu2018ano_pred ; ramachandra2020street that torwards real-world introduced more complex scenes with crowded traffic, yet remained unimodal, fixed-view, and emphasized foreground representation. Larger-scale datasets sultani2018real ; wu2020not ; zhu2024advancing ) improved diversity and duration, but still relied on ground-view perspectives and coarse anomaly labels, lacking spatial localization or causal reasoning. Methodologically, the field evolved from handcrafted features to learned representations adam2008robust ; zhao2011online ; li2013anomaly and behavior modeling mehran2009abnormal ; li2018resound ; kim2009observe . Recent approaches span reconstruction/prediction-based hasan2016learning ; lv2023unbiased ; gong2019memorizing ; xu2017detecting ; yang2023video ; liu2018future , object-centric liu2021hybrid ; li2021variational ; lv2021learning ; tian2022pixel ; gao2023atta ; zhao2024segment , distribution-aware hirschorn2023normalizing ; tu2024self ; micorek2024mulde ; leng2024beyond , and llm-driven paradigms du2024uncovering ; zanella2024harnessing ; lv2024video ; zhang2024holmes , with growing emphasis on generalization and scene dependency. However, existing benchmarks and methods largely overlook aweial-specific challenges such as extreme viewpoint shifts and scale variation. Moreover, the absence of multimodal and reasoning-oriented annotations limits fine-grained analysis.
Aerial Anomaly Understanding. Anomaly detection in aerial videos remains underexplored. Early pioneer datasets bonetto2015privacy ; bozcan2021context ; jin2022anomaly ; tran2023uit adopt aerial perspectives but provide only frame-level or coarse labels, limiting fine-grained analysis. Existing methods focus on motion cues—optical flow bozcan2021context , reconstruction-based schemesle2025hstforu , or spatiotemporal modeling with 3D CNNs tran2023uit and Transformer jin2022anomaly , but rarely support precise region-level reasoning. While multimodal large language models (MLLMs) achiam2023gpt ; hu2024minicpm ; chen2024internvl ; chen2024far ; li2024llavanext-strong ; yang2024qwen2 have advanced semantic understanding in ground-view tasks, their application to aerial scenarios is limited. Current approaches bharadwaj2024vanebench ; tang2024hawk ; zhang2024holmes often lack explicit, grounded reasoning and rely on post hoc explanations. To bridge this gap, we introduce a reasoning-centric aerial anomaly dataset with detailed spatial-temporal annotations and dynamic reasoning paths.
@@ 3 The A2Seek Dataset
Existing video anomaly detection methods primarily rely on fixed-view ground-based cameras, which are designed for limited fields of view and relatively static backgrounds. These methods face significant limitations when applied to drone-view videos, which involve frequent viewpoint changes, scale variations, dynamic occlusions, and complex environmental disturbances (e.g., lighting changes, weather variations). These factors significantly increase the challenges of spatial localization and semantic generalization anomaection. To address these challenges, we introduce A2Seek, a reasoning-centric aerial anomaly understandunderstanding benchmark. It supports precise spatial localization of anomalies (“Where is the anomaly?”) and in-depth semantic reasoning explanations (“Why is it anomalous?”). The dataset spans diverse real-world scenarios and anomaly types, providing high-resolution RGB and infrared video data with detailed frame-level labels, region-level annotations, and structured natural language reasoning explanations. Figure 2 illustrates the dataset’s diversity, showcasing various scenes, altitudes, speeds, weather conditions, and times of day. Unlike traditional ground-based datasets, which focus on static viewpoints and limited environmental variations, A2Seek leverages the dynamic and expansive nature of UAV perspectives, making it a more challenging and realistic benchmark. This benchmark aims to advance research on generalization, robustness, and interpretability in aerial anomaly understanding.
@@ 3.1 Data Collection and Annotation
The A2Seek dataset was collected using a DJI M30T drone equipped with wide-angle, telephoto, and infrared cameras, operating at varying altitudes (10–60 meters) and speeds (0–20 m/s) to capture diverse scenarios. Flight trajectories included hovering, linear cruising, curved circling, and area scanning, reflecting dynamic viewpoint changes. The dataset comprises 542 untrimmed 4K videos (over 23 hours) recorded across 10 campus environments, featuring over 20 anomaly categories (e.g., falling, fighting, jaywalking) under various conditions (day/night, clear/foggy, etc.). To ensure high-quality and comprehensive labeling, A2Seek employs a rigorous multi-level annotation framework. This process includes spatiotemporal bounding boxes, fine-grained categories, and structured reasoning paths. A multi-stage annotation process was designed, encompassing trigger identification, anomaly diagnosis, causal reasoning, reflection, and seeking phases (as shown in Figure 1e). Specific formatting rules were applied to align model outputs with human annotations, ensuring consistency and interpretability. During this process, the model generated multiple responses, from which professional annotators selected and refined the most appropriate ones. The finalized annotations were integrated into a reasoning-centric framework, supporting precise anomaly detection and facilitating in-depth semantic reasoning.
@@ 3.2 Dataset Characteristics and Comparative Analysis
Table 1: Comparison across our proposed A2Seek and existing video anomaly detection datasets.
Perspective Dataset Frames Scene Count Anomaly Types Resolution Scene Dependency Scale Variation Reasoning Annotation Multi- modal
@@
Surveillance CUHK Avenuelu2013abnormal 30,652 26,832 3,820 1 5 640$\times$360
Shanghai Techluo2017revisit 317,398 300,308 17,090 13 11 856$\times$480
Street Sceneramachandra2020street 203,257 159,341 43,916 1 17 1,280$\times$720
Subway Entranceadam2008robust 144,250 132,138 12,112 1 5 512$\times$384
Subway Exitadam2008robust 64,901 60,410 4,491 1 3 512$\times$384
IITB Corridorchan2017anticipating 483,566 301,999 181,567 1 10 1,920$\times$1,080
UCSD Ped1li2013anomaly 14,000 9,995 4,005 1 5 238$\times$158
UCSD Ped2li2013anomaly 4,560 2,924 1,636 1 5 360$\times$240
UMNraghavendra2006unusual 7,741 6,165 1,576 3 1 320$\times$240
UBnormalacsintoae2022ubnormal 236,902 147,887 89,015 29 22 1,080$\times$720
NWPU Campuscao2023new 1,466,073 1,400,807 65,266 43 28 2,048$\times$1,536
MSADzhu2024advancing 447,236 308,792 138,444 14 55 1920$\times$1080
Drone-view Minidronebonetto2015privacy 23,295 14,821 8,474 1 10 1,280$\times$720
AU-AIR-Anomalybozcan2021context 32,823 30,000 2,823 1 8 1,920$\times$1,080
Drone-Anomalyjin2022anomaly 87,488 51,635 35,853 7 1 640$\times$640
UIT-ADronetran2023uit 206,194 142,709 63,485 3 10 1920$\times$1080
A2Seek (Ours) 2,485,859 2,087,160 398,699 27 20 3,840$\times$2,160
Table 1 compares A2Seek with existing aerial and ground-based anomaly detection datasets. As the largest aerial anomaly detection dataset, A2Seek comprises over 2.4 million frames, including 398k frames containing anomalies, significantly surpassing other drone-view datasets in both scale and diversity. The dataset spans 27 distinct sub-scenes, such as roadways, sidewalks, playgrounds, and industrial zones, providing extensive coverage of real-world scenarios. Beyond its scale and diversity, A2Seek is the first aerial anomaly detection dataset to provide natural language reasoning annotations with precise spatial localization labels. These annotations enable models to perform deep anomaly understanding and reasoning, bridging the gap between detection and interpretability. Furthermore, A2Seek explicitly addresses challenges unique to aerial perspectives, such as scene dependency, subtle anomalies, complex backgrounds, and scale variations, while also integrating multimodal data (RGB and infrared) to handle diverse environmental conditions. In summary, A2Seek excels in data modalities, scene richness, and fine-grained annotations, directly addressing the two core challenges of aerial anomaly detection: spatial localization and semantic generalization. By establishing a robust data foundation and offering clear research directions, A2Seek sets a new benchmark for advancing aerial video anomaly detection and promoting deeper exploration in this domain. More details about our A2Seek data are reported in Appendix A.

### v2

@@ Abstract
While unmanned aerial vehicles (UAVs) offer wide-area, high-altitude coverage for anomaly detection, they face challenges such as dynamic viewpoints, scale variations, and complex scenes. Existing datasets and methods, mainly designed for fixed ground-level views, struggle to adapt to these conditions, leading to significant performance drops in drone-view scenarios. To bridge this gap, we introduce A2Seek (Aerial Anomaly Seek), a large-scale, reasoning-centric benchmark dataset for aerial anomaly understanding. This dataset covers various scenarios and environmental conditions, providing high-resolution real-world aerial videos with detailed annotations, including anomaly categories, frame-level timestamps, region-level bounding boxes, and natural language explanations for causal reasoning. Building on this dataset, we propose A2Seek-R1, a novel reasoning framework that generalizes R1-style strategies to aerial anomaly understanding, enabling a deeper understanding of “Where” anomalies occur and “Why” they happen in aerial frames. To this end, A2Seek-R1 first employs a graph-of-thought (GoT)-guided supervised fine-tuning approach to activate the model’s latent reasoning capabilities on A2Seek. Then, we introduce Aerial Group Relative Policy Optimization (A-GRPO) to design rule-based reward functions tailored to aerial scenarios. Furthermore, we propose a novel “seeking” mechanism that simulates UAV flight behavior by directing the model’s attention to informative regions. Extensive experiments demonstrate that A2Seek-R1 achieves up to a 22.04% improvement in AP for prediction accuracy and a 13.9% gain in mIoU for anomaly localization, exhibiting strong generalization across complex environments and out-of-distribution scenarios. Our dataset and code are released at https://2-mo.github.io/A2Seek/.
@@ 1 Introduction
Traditional anomaly detection lu2013abnormal; liu2018ano_pred; cao2023new relies on fixed-view cameras and primarily focuses on anomaly classification, offering limited semantic interpretation. Their static perspectives and narrow fields of view significantly limit their effectiveness in monitoring large and dynamic environments zhu2024advancing. With the rapid advancement of unmanned aerial vehicle (UAV) technology, aerial surveillance has emerged as a powerful paradigm for wide-area anomaly detection. Drone‐view footage introduces frequent viewpoint shifts, scale changes, complex backgrounds and occlusions, as well as environmental disturbances (lighting, weather, moving shadows) cao2021visdrone; du2018unmanned. Crucially, anomalous regions in aerial scenes are often subtle, spatially sparse, and occupy only a small portion of the field of view, making them difficult to perceive. Even when alarms are triggered, human observers often struggle to localize these subtle, spatially sparse events. Moreover, traditional models rely on stable spatial priors invalidated when a drone’s pose constantly changes. Consequently, ground‐view approaches fail to generalize to aerial data, exhibiting poor robustness to dynamic scenes and variable spatial distributions. As Figure 1 illustrates, anomaly detection in drone-view footage requires precise spatial localization (“Where is the anomaly?”) and semantic interpretation (“Why is it an anomaly?”). Drone-captured videos often span large, dynamic environments, where subtle anomalies can be easily overlooked or misinterpreted without robust contextual reasoning. These factors demand models capable of adaptively focusing on critical regions and abstracting high-level reasoning to explain anomalies. In the following, we distill these into two core problems critical for robust aerial anomaly understanding.
@@
CHALLENGE 1: Spatial Localization in Wide-Area, Dynamic Scenes. Traditional ground-based anomaly detection typically relies on fixed-view cameras, where anomalies are often prominent and relatively easy to identify. However, aerial anomaly detection is fundamentally different. Drone-view videos cover expansive and dynamic scenes with complex backgrounds and frequent motion. Anomalies are often small, sparse, and easily obscured by cluttered surroundings, making them difficult to detect even when anomalies are flagged (e.g., a person falling in Figure 1a). Therefore, aerial anomaly detection requires not only determining whether an event is anomalous but also providing accurate localization of the anomalous objects to support anomaly judgment.
@@
Building on this benchmark, we propose A2Seek-R1, a novel reinforcement fine-tuning framework designed to enhance the reasoning capabilities of models for aerial anomaly understanding. A2Seek-R1 first employs a graph-of-thought (GoT)-guided supervised fine-tuning (SFT) approach, which activates the model’s latent reasoning capabilities by leveraging structured reasoning annotations in the A2Seek dataset. These annotations consist of optional stages, including trigger, diagnosis, reasoning, reflection, and seeking, effectively guiding the model to handle anomalies of varying complexity in a progressive manner. Among them, seeking is set as a potential region of interest for the model in video frames with insufficient information, such as blurry or occluded images, thus achieving a new type of seeking mechanism that simulates the flight behavior of unmanned aerial vehicles, enabling the model to dynamically focus on specific regions of interest. Second, it introduces a tailored extension of Group Relative Policy Optimization (GRPO), termed A-GRPO, specifically designed for aerial anomaly understanding. A-GRPO extends the original accuracy and format function rewards by incorporating localization and seeking rewards. Localization rewards enhance the model’s spatial understanding of anomaly regions, while seeking rewards focus on aligning the model’s predictions with human annotations of anomaly candidate areas, ensuring the extraction of valuable spatial information for better understanding. Additionally, to address the diverse perspectives of drones, a length reward function is introduced to encourage concise responses in simple scenarios and allocate more computational effort to complex situations. By combining these components, A2Seek-R1 achieves precise spatial localization and robust reasoning for aerial anomalies, setting a new benchmark for anomaly understanding in complex, real-world environments.
@@ 2 Related Work
Video Anomaly Detection. Early efforts focused on single-scene datasets li2013anomaly; lu2013abnormal using fixed-view RGB cameras for pedestrian anomaly detection. Later datasets liu2018ano_pred; ramachandra2020street; degardin2021iterative; lv2021localizing that towards real-world introduced more complex scenes with crowded traffic, yet remained unimodal, fixed-view, and emphasized foreground representation. Larger-scale datasets sultani2018real; wu2020not; wan2021anomaly; NEURIPS2024_a3c5af1f improved diversity and duration, but still relied on ground-view perspectives and coarse anomaly labels, lacking spatial localization or causal reasoning. Methodologically, the field evolved from handcrafted features to learned representations adam2008robust; zhao2011online; li2013anomaly and behavior modeling mehran2009abnormal; li2018resound; kim2009observe. Recent approaches span reconstruction/prediction-based hasan2016learning; lv2023unbiased; gong2019memorizing; xu2017detecting; yang2023video; liu2018future, object-centric liu2021hybrid; li2021variational; lv2021learning; tian2022pixel; gao2023atta; zhao2024segment, distribution-aware hirschorn2023normalizing; tu2024self; micorek2024mulde; leng2024beyond, and llm-driven paradigms du2024uncovering; zanella2024harnessing; lv2024video; zhang2024holmes, with growing emphasis on generalization and scene dependency. Recent methods zhu2025vaur1; huang2025vadr1 incorporate chain-of-reasoning mechanisms to enhance anomaly understanding, enabling more interpretable and goal-directed decision-making. However, existing benchmarks and methods largely overlook aerial-specific challenges such as extreme viewpoint shifts and scale variation. Moreover, the absence of multimodal and reasoning-oriented annotations limits fine-grained analysis.
Aerial Anomaly Understanding. Anomaly detection in aerial videos remains underexplored. Early pioneer datasets bonetto2015privacy; bozcan2021context; jin2022anomaly; tran2023uit adopt aerial perspectives but provide only frame-level or coarse labels, limiting fine-grained analysis. Existing methods focus on motion cues, including optical flow bozcan2021context, reconstruction-based schemesle2025hstforu, or spatiotemporal modeling with 3D CNNs tran2023uit and Transformer jin2022anomaly, but rarely support precise region-level reasoning. While multimodal large language models (MLLMs) achiam2023gpt; hu2024minicpm; chen2024internvl; chen2024far; li2024llavanext-strong; yang2024qwen2 have advanced semantic understanding in ground-view tasks, their application to aerial scenarios is limited. Current approaches bharadwaj2024vanebench; tang2024hawk; zhang2024holmes often lack explicit, grounded reasoning and rely on post hoc explanations. To bridge this gap, we introduce a reasoning-centric aerial anomaly dataset with fine-grained spatial-temporal annotations and dynamic reasoning trajectories.
@@ 3 The A2Seek Dataset
Existing video anomaly detection methods primarily rely on fixed-view ground-based cameras, which are designed for limited fields of view and relatively static backgrounds. These methods face significant limitations when applied to drone-view videos, which involve frequent viewpoint changes, scale variations, dynamic occlusions, and complex environmental disturbances (e.g., lighting changes, weather variations). These factors significantly increase the challenges of spatial localization and semantic generalization of anomaly detection. To address these challenges, we introduce A2Seek, a reasoning-centric aerial anomaly understanding benchmark. It supports precise spatial localization of anomalies (“Where is the anomaly?”) and in-depth semantic reasoning explanations (“Why is it anomalous?”). The dataset spans diverse real-world scenarios and anomaly types, providing high-resolution RGB and infrared video data with detailed frame-level labels, region-level annotations, and structured natural language reasoning explanations. Figure 2 illustrates the dataset’s diversity, showcasing various scenes, altitudes, speeds, weather conditions, and times of day. Unlike traditional ground-based datasets, which focus on static viewpoints and limited environmental variations, A2Seek leverages the dynamic and expansive nature of UAV perspectives, making it a more challenging and realistic benchmark. This benchmark aims to advance research on generalization, robustness, and interpretability in aerial anomaly understanding.
@@ 3.1 Data Collection and Annotation
The A2Seek dataset was collected using a DJI M30T drone equipped with wide-angle, telephoto, and infrared cameras. Flights were conducted at varying altitudes (10 to 60 meters) and speeds (0 to 20 m/s) to capture diverse aerial perspectives. Trajectory patterns included hovering, linear cruising, curved circling, and area scanning, enabling dynamic viewpoint shifts across scenes. In total, A2Seek comprises 542 untrimmed 4K videos (over 23 hours), recorded across 10 campus environments, covering more than 20 types of anomalous events (e.g., falling, fighting, jaywalking) under varying conditions such as day/night, clear/foggy weather, and so on.
The anomaly categories in A2Seek were carefully curated based on the principle of “potential disruption to campus public safety or order,” rather than broadly labeling daily activities as abnormal. For instance, running on a playground is not considered anomalous, whereas running in an academic building area is. Similarly, riding a bicycle on the road is normal, but riding on sidewalks is considered anomalous. Drawing from classic VAD datasets li2013anomaly; lu2013abnormal; luo2017revisit; cao2023new, we focus on behavior categories with strong relevance to real-world campus safety. Anomalies are categorized into three risk levels. High-risk events such as carrying weapons, fighting, and robbery require immediate intervention. Medium-risk events such as running, falling, animal intrusions, and illegal cycling may escalate but do not cause immediate harm. Low-risk events such as loitering, littering, and walking on grass pertain to order maintenance or environmental management.
To ensure high-quality and comprehensive labeling, A2Seek employs a rigorous multi-level annotation framework. This process includes spatiotemporal bounding boxes, fine-grained categories, and structured reasoning paths. A multi-stage annotation process was designed, encompassing trigger identification, anomaly diagnosis, causal reasoning, reflection, and seeking phases (as shown in Figure 1e). Specific formatting rules were enforced to align model-generated responses with human annotations. During annotation, the model produced multiple candidate explanations, from which professional annotators selected and refined the most appropriate. The final annotations are integrated into a reasoning-centric framework, enabling both precise anomaly localization and high-level semantic understanding. For privacy, all facial regions and identity-related visual attributes are blurred. Further procedural details are provided in Appendix D.
@@ 3.2 Dataset Characteristics and Comparative Analysis
Table 1: Comparison of A2Seek with existing video anomaly detection datasets ( denotes web-sourced datasets; denotes simulated or virtual datasets).
Perspective Dataset Frames
Scene
Count
Anomaly
Types
Resolution
Scene
Dependency
Scale
Variation
Reasoning
Annotation
Multi-
modal
@@
[…]
Drone-Anomaly jin2022anomaly 87,488 51,635 35,853 7 1 $\times640$
[…]
Table 1 compares A2Seek with existing aerial and ground-based anomaly detection datasets. As the largest aerial anomaly detection dataset, A2Seek comprises over 2.4 million frames, including 398k frames containing anomalies, significantly surpassing other drone-view datasets in both scale and diversity. The dataset covers 10 major scene categories, such as roadways, sidewalks, playgrounds, and industrial zones. These categories are further divided into 27 specific sub-scenes, providing comprehensive coverage of complex real-world environments.
Beyond its scale and diversity, A2Seek is the first aerial anomaly detection dataset to provide natural language reasoning annotations with precise spatial localization labels. These annotations enable models to perform deep anomaly understanding and reasoning, bridging the gap between detection and interpretability. Furthermore, A2Seek explicitly addresses challenges unique to aerial perspectives, such as scene dependency, subtle anomalies, complex backgrounds, and scale variations, while also integrating multimodal data (RGB and infrared) to handle diverse environmental conditions. In summary, A2Seek excels in data modalities, scene richness, and fine-grained annotations, directly addressing the two core challenges of aerial anomaly detection: spatial localization and semantic generalization. By establishing a robust data foundation and offering clear research directions, A2Seek sets a new benchmark for advancing aerial video anomaly detection and promoting deeper exploration in this domain. More details about our A2Seek data are reported in Appendix A.